# Rank-1 Approximation of Inverse Fisher for Natural Policy Gradients in Deep Reinforcement Learning

### v1

@@ 1 Introduction
One approach to reduce this sub-optimality and achieve stable training with large steps is to incorporate the geometry of the loss manifold into account and thus consider natural gradient descent methods (Amari, 1998). Specifically, when applied to policy gradient RL, where policies are modeled as probability distributions, the Fisher information matrix (FIM) is used to precondition the policy gradients, which gives rise to natural policy gradient (NPG) methods (Kakade, 2001; Bagnell and Schneider, 2003; Peters et al., 2005). NPG algorithms are of interest because they generally tend to converge faster and have superior performance compared to their standard policy gradient counterparts (Grondman et al., 2012). More recently, NPGs established the roots of some of the most popular RL algorithms used in practice, including Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a) and Proximal Policy Optimization (PPO) (Schulman et al., 2017). Furthermore, the study of NPG methods is also motivated by their success in real-world applications (Richter et al., 2006; Su et al., 2019).
@@
The goal of this work is to bridge the theoretical convergence properties of NPG with a practical, fast, and computationally feasible approach to invert FIM for high-dimensional reinforcement learning tasks. In this work, we use the empirical FIM as an alternative to the FIM and incorporate the Sherman-Morrison formula for efficient inversions. We propose a rank-1 approximation to avoid explicit matrix operations, keeping the time complexity at $\mathcal{O}(d)$ without the need for inner loops like conjugate gradient, thus balancing computational cost and accuracy of curvature information.
@@
    We provide a fast and efficient way to compute the inverse FIM, which can be computed in $\mathcal{O}(d)$ complexity rather than $\mathcal{O}(d^{3})$ for naive inversion of FIM, $d$ denotes the dimension of the policy parameter vector $\theta$.
@@
    We propose a novel policy gradient algorithm using this update: SM-ActorCritic (SMAC). We show that SMAC achieves faster convergence in diverse OpenAI Gym tasks (classic control and MuJoCo environments), compared to vanilla policy gradient, Adam, and trust-region methods that use Conjugate Gradient solvers to approximate the direction of natural policy gradient.
@@ 2 Related Work
Gradient descent updates all the parameters of a model equally, meaning that weights are considered to be inside a Euclidean space. However, the loss function of the model actually induces a Riemannian manifold space. Therefore, the size of a gradient step does not reflect the true scale of change of the model’s loss. Consequently, standard gradient descent methods are highly sensitive to the choice of step size. To avoid this issue, Amari first framed gradient-based learning on a Riemannian manifold, introducing the natural gradient descent algorithm, which uses the update direction $F(\theta)^{-1}\nabla J(\theta)$, with the FIM $F(\theta)$ Amari (1998). Here, the FIM is a covariant measure of the amount of information contained in the parameters Amari (1998). Kakade extended this idea to reinforcement learning and coined the natural policy gradient (NPG) algorithm Kakade (2001). However, the computational cost associated with calculating the FIM, especially in models with a large number of parameters, can be prohibitively expensive. Therefore, practical implementations of natural gradients rely on approximations instead. We thus summarize the related work on these approximations as follows.
@@
Hessian‑free optimization Martens (2010), another method for computing natural gradients, has seen widespread use in reinforcement learning, particularly due to its superior performance in TRPO Schulman et al. (2015a). This approach employs the conjugate gradient (CG) method to solve the associated linear system, thereby avoiding explicitly storing FIM, keeping memory usage comparable to a single forward–backward pass. However, CG requires several additional forward–backward passes, which reduces computational efficiency. In addition, due to the limited precision of CG estimation, practitioners usually perform an extra KL-based line search to adapt the step size (Schulman et al., 2015a).
@@
The empirical Fisher (EF) estimates the Fisher information matrix by replacing the exact expectation over gradient outer products with the outer product of the gradients computed from individual samples Yang et al. (2022). However, the empirical Fisher’s simple estimate has been frequently criticized for deviating significantly from the true Fisher, and its limitations have been analyzed in many studies Martens (2020); Kunstner et al. (2019). Nevertheless, its very low computational cost has enabled a broad range of successful applications Roux et al. (2007); Ren and Goldfarb (2019); Wu et al. (2024). Compared with the extensive use of K‑FAC and Hessian‑free methods in RL, the EF approach remains under‑explored and is seldom applied in this domain.
@@ 3 Preliminaries
We consider an episodic MDP $(\mathcal{S},\mathcal{A},P,R,\gamma)$ with a continuous or discrete state space $\mathcal{S}$, action space $\mathcal{A}$, transition kernel $P(s^{\prime}|s,a)$, bounded reward function $r:S\times A\mapsto[-R,R]$, and discount factor $\gamma\in(0,1)$. We also consider a stochastic policy $\pi_{\theta}(a|s)$ with parameters $\theta\in\mathbb{R}^{d}$ that induces a trajectory distribution $\pi_{\theta}(\tau_{i})$ over $\tau_{i}=(s_{0}^{i},a_{0}^{i},s_{1}^{i},a_{1}^{i},\dots)$.
@@
To maximize Eq. 1, the advantage actor-critic (A2C) (Mnih et al., 2016) algorithm uses an advantage function $A^{\pi_{\theta}}(s,a)=Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)$, defined through the critic’s state-value and action-value functions
$V^{\pi_{\theta}}(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[G_{t}\mid s_{t}=s\right],\qquad Q^{\pi_{\theta}}(s,a)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[G_{t}\mid s_{t}=s,a_{t}=a\right],$
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\Bigl[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\;A^{\pi_{\theta}}(s_{t},a_{t})\Bigr].$
@@
For the exponential‐family policy $\pi_{\theta}(\cdot\mid s)$, the natural Riemannian metric in parameter space is the Fisher information matrix
$\bm{F}_{s}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\!\bigl[\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\!\top}\bigr].$ (2)
For a preconditioning matrix $\bm{F}_{\rho}(\theta)$, this leads to the natural gradient update
@@
where $d_{\rho}^{\pi_{\theta}}$ represents the state visitation distribution under a policy $\pi_{\theta}$ and initial state distribution $\rho$:
@@
Replacing the expectation in ${\bm{F}}(\theta)$ with a sample average over a batch $\{\tau_{i}\}_{i=1}^{N}$ gives the empirical Fisher approximation
$\hat{\bm{F}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}.$
@@
The Sherman–Morrison formula efficiently computes the inverse of a matrix after a “rank-1 update” has been applied to it, provided that the inverse of the original matrix was already known Sherman and Morrison (1950). Suppose $\bm{X}\in\mathbb{R}^{n\times n}$ is an invertible square matrix and $\bm{u},\bm{v}\in\mathbb{R}^{n}$ are column vectors. Then $\bm{X}+\bm{u}\bm{v}^{\top}$ is invertible iff
@@
If $\bm{X}^{-1}$ is already known, this matrix inversion formula can bypass computing the inverse $({\bm{X}}+{\bm{u}}{\bm{v}}^{T})^{-1}$ directly, avoiding the cubic scaling with the dimension of $\bm{X}$}. Instead, the update only requires three vector operations, which scale linearly with the dimension of $\bm{X}$}. This update can use the full power of current GPU hardware platforms, since computing vector operations can be heavily parallelized. At the same time, it avoids the time-intensive matrix inversion processes. Specifically, methods like Singular Value Decomposition (SVD), which decomposes a matrix into its singular values and vectors, and LU decomposition, which factors a matrix into a lower triangular matrix (L) and an upper triangular matrix (U), are much slower than the simple operations used in our approach.
@@ 4 NPG Approximation
We adopt a matrix-free scheme for approximating the FIM to reduce memory and computational cost. At each step, we form a local empirical Fisher from the current batch, add a damping term $\lambda$, and apply the Sherman–Morrison formula to compute the inverse–vector product ${\bm{F}}(\theta)^{-1}g^{k}$ directly. This approach requires $\mathcal{O}(d)$ memory.
@@ 4.1 Matrix-Free Update
$({\hat{\bm{F}}^{k}})^{-1}g^{k}$ $=(\hat{\bm{F}}^{k-1}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top})^{-1}g^{k}$
$\Delta\theta=\eta(\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top})^{-1}\log\pi_{\theta}(a|s).$
However, the time and space complexity of explicitly calculating and storing the FIM is $\mathcal{O}(d^{2})$, which is infeasible for large-scale models. In this work, we instead use the Sherman–Morrison formula to directly compute the product of the inverse empirical Fisher and the policy gradient.
We consider a one-sample empirical Fisher with damping:
$\hat{\bm{F}}^{k}=\lambda\bm{I}+\bm{F}^{k}(\theta)=\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}$
where $\lambda>0$ is a fixed damping coefficient.
Applying the Sherman–Morrison formula yields:
$({\hat{\bm{F}}^{k}})^{-1}$ $=(\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top})^{-1}$
$=\frac{1}{\lambda}\bm{I}-\frac{\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}}{\lambda^{2}+\lambda\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)}$
$=\frac{1}{\lambda}\bm{I}-\frac{\bm{F}^{k}(\theta)}{\lambda^{2}+\lambda Tr(\bm{F}^{k}(\theta))}.$ (4)
Multiplying the inverse Fisher with the gradient gives the natural policy gradient update direction:
$\Delta\theta_{k}=\eta\,({\hat{\bm{F}}^{k}})^{-1}g^{k}=\eta\left[\frac{1}{\lambda}g^{k}-\frac{\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}g^{k}}{\lambda^{2}+\lambda\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)}\right]$ (5)
where $g^{k}=\nabla_{\theta}\log\pi_{\theta}(a|s)A(s,a)$ is the one-sample policy gradient update direction, $A(s,a)$ is the advantage estimate and $\eta$ is the step size.
@@ 4.2 Convergence Analysis
We take inspiration from the global convergence guarantees provided by Agarwal et al. (2021) and a follow-up work by Liu et al. (2020), which improved the global convergence results. To prove the global convergence, we take a standard Lipschitz assumption and take some previously proved bounds of the norm of exact policy gradient and the policy gradient defined till horizon $H$ from Xu et al. (2020). We provide a theoretical analysis showing that the Sherman–Morrison rank-1 approximation to the inverse Fisher achieves global convergence. Under strong convexity assumptions (already taken in Kakade (2001), Peters et al. (2005)) of the FIM, we show that the average performance between the Sherman-Morrison update and the optimal policy $\pi^{*}$ can be bounded. We defer the proof to the appendix, and state the final theorem here.
@@
In total, stochastic PG samples $\mathcal{O}\left(\frac{\sigma^{2}}{(1-\gamma)^{2}\varepsilon^{4}}\right)$ trajectories, where $\varepsilon_{\text{bias}}$ is the approximation error from the advantage function, and $\varepsilon$ is an arbitrary precision level.
@@
Our proof relies on the fact that performance bound given by Lemma 4 Liu et al. (2020) and we prove that this result holds with a sample complexity of $KN=\mathcal{O}(\frac{\sigma^{2}}{(1-\gamma)^{2}\varepsilon^{4}})$, where $N$ is the number of trajectories per update and $K$ is the number of updates. This shows that to get an $\varepsilon$ accurate policy, one needs $\mathcal{O}(\frac{1}{\varepsilon^{2}})$ samples, and when the horizon $H$ and number of samples $N$ are large enough, then $\sigma$ can be reduced to arbitrary precision. Detailed derivations can be found in Appendix 1.

### v2

@@ 1 Introduction
One approach to reduce this sub-optimality and achieve stable training with large steps is to incorporate the geometry of the loss manifold into account and thus consider natural gradients (Amari, 1998). Specifically, when applied to policy gradient RL, where policies are modeled as probability distributions, the Fisher information matrix (FIM) is used to precondition the policy gradients, which gives rise to natural policy gradient (NPG) methods (Kakade, 2001; Bagnell and Schneider, 2003; Peters et al., 2005). NPG algorithms are of interest because they generally tend to converge faster and have superior performance compared to their standard policy gradient counterparts (Grondman et al., 2012). More recently, NPGs established the roots of some of the most popular RL algorithms used in practice, including Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a), Proximal Policy Optimization (PPO) (Schulman et al., 2017). Furthermore, the study of NPG methods is also motivated by their success in real-world applications (Richter et al., 2006; Su et al., 2019).
@@
The goal of this work is to bridge the theoretical convergence properties of NPG with a practical, fast, and computationally feasible approach to invert FIM for high-dimensional reinforcement learning tasks. In this work, we use the empirical FIM as an alternative to the FIM and incorporate the Sherman-Morrison formula for efficient inversions. We propose a rank-1 approximation to avoid explicit matrix operations, keeping the time complexity at $\mathcal{O}(d)$ for a policy parameterized by a vector $\theta$ of dimension $d$, without the need for inner loops like conjugate gradient, thus balancing computational cost and accuracy of curvature information.
@@
    We provide a fast and efficient way to compute the inverse FIM, which can be computed in $\mathcal{O}(d)$ complexity rather than $\mathcal{O}(d^{3})$ for naive inversion of FIM.
@@
    We propose a novel policy gradient algorithm using this update: Sherman-Morrison Actor Critic (SMAC). We show that SMAC achieves faster convergence in diverse OpenAI Gym tasks (classic control and MuJoCo environments (Todorov et al., 2012)), compared to vanilla policy gradient, Adam, and trust-region methods that use Conjugate Gradient solvers to approximate the direction of natural policy gradient.
@@ 2 Related Work
Gradient descent updates all the parameters of a model equally, meaning that weights are considered to be inside a Euclidean space. However, the loss function of the model actually induces a Riemannian manifold space. Therefore, the size of a gradient step does not reflect the true scale of change of the model’s loss. Consequently, standard gradient descent methods are highly sensitive to the choice of step size. To avoid this issue, Amari first framed gradient-based learning on a Riemannian manifold, introducing the natural gradient descent algorithm, which uses the update direction $F(\theta)^{-1}\nabla J(\theta)$, with the FIM $F(\theta)$ and objective $J(\theta)$ (Amari, 1998). Here, the FIM is a covariant measure of the amount of information contained in the parameters (Amari, 1998). Kakade extended this idea to reinforcement learning and coined the natural policy gradient (NPG) algorithm (Kakade, 2001). However, the computational cost associated with calculating the FIM, especially in models with a large number of parameters, can be prohibitively expensive. Therefore, practical implementations of natural gradients rely on approximations instead. We thus summarize the related work on these approximations as follows.
@@
Hessian‑free optimization (Martens, 2010), another method for computing natural gradients, has seen widespread use in reinforcement learning, particularly due to its superior performance in TRPO (Schulman et al., 2015a). This approach employs the conjugate gradient (CG) method to solve the associated linear system, thereby avoiding explicitly storing FIM, keeping memory usage comparable to a single forward–backward pass. However, CG requires several additional forward–backward passes, which reduces computational efficiency. In addition, due to the limited precision of CG estimation, practitioners usually perform an extra KL-based line search to adapt the step size (Schulman et al., 2015a).
@@
The empirical Fisher (EF) estimates the Fisher information matrix by replacing the exact expectation over gradient outer products with the outer product of the gradients computed from individual samples (Yang et al., 2022). However, the empirical Fisher’s simple estimate has been frequently criticized for deviating significantly from the true Fisher, and its limitations have been analyzed in many studies (Martens, 2020; Kunstner et al., 2019). Nevertheless, its very low computational cost has enabled a broad range of successful applications (Roux et al., 2007; Ren and Goldfarb, 2019; Wu et al., 2024). Compared with the extensive use of K‑FAC and Hessian‑free methods in RL, the EF approach remains under‑explored and is seldom applied in this domain.
@@ 3 Preliminaries
We consider an MDP $(\mathcal{S},\mathcal{A},P,R,\gamma)$ with a continuous or discrete state space $\mathcal{S}$, action space $\mathcal{A}$, transition kernel $P(s^{\prime}|s,a)$, bounded reward function $r:S\times A\mapsto[-R,R]$, and discount factor $\gamma\in(0,1)$. We also consider a stochastic policy $\pi_{\theta}(a|s)$ with parameters $\theta\in\mathbb{R}^{d}$ that induces a trajectory distribution $\pi_{\theta}(\tau_{i})$ over $\tau_{i}=(s_{0}^{i},a_{0}^{i},s_{1}^{i},a_{1}^{i},\dots)$.
@@
To maximize Eq. 1, the advantage actor-critic (A2C) (Mnih et al., 2016) algorithm uses an advantage function A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s), defined through the critic’s state-value and action-value functions
$V(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[G_{t}\mid s_{t}=s\right],\qquad Q(s,a)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[G_{t}\mid s_{t}=s,a_{t}=a\right],$
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\Bigl[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\;A(s_{t},a_{t})\Bigr].$
@@
For the exponential‐family policy $\pi_{\theta}(\cdot\mid s)$, the natural Riemannian metric in parameter space can be expressed in terms of state visitation distribution induced by the same policy $\pi_{\theta}(\cdot\mid s)$ as
$\bm{F}_{s}(\theta)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\!\bigl[\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\!\top}\bigr].$ (2)
The invariant FIM $\bm{F}_{\rho}(\theta)$ can be obtained by averaging over the state visitation distribution leading to the natural gradient update
@@
where $\eta$ is the step size and $d_{\rho}^{\pi_{\theta}}$ represents the state visitation distribution under a policy $\pi_{\theta}$ and initial state distribution $\rho$:
@@
Replacing the expectation in ${\bm{F}}(\theta)$ with a sample average over one single trajectory $\tau=\{(s_{i},a_{i})\}_{t=0}^{T-1}$ gives the empirical Fisher approximation
$\hat{\bm{F}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log\pi_{\theta}(a_{i}|s_{i})\nabla_{\theta}\log\pi_{\theta}(a_{i}|s_{i})^{\top}.$
@@
The Sherman–Morrison formula efficiently computes the inverse of a matrix after a “rank-1 update” has been applied to it, provided that the inverse of the original matrix was already known (Sherman and Morrison, 1950). Suppose $\bm{X}\in\mathbb{R}^{n\times n}$ is an invertible square matrix and $\bm{u},\bm{v}\in\mathbb{R}^{n}$ are column vectors. Then $\bm{X}+\bm{u}\bm{v}^{\top}$ is invertible iff
@@
If $\bm{X}^{-1}$ is already known, this matrix inversion formula can bypass computing the inverse $({\bm{X}}+{\bm{u}}{\bm{v}}^{T})^{-1}$ directly, avoiding the cubic scaling with the dimension of $\bm{X}$}. Instead, the update only requires three vector operations, which scale linearly with the dimension of $\bm{X}$}. This update can use the full power of current GPU hardware platforms, since computing vector operations can be heavily parallelized. At the same time, it avoids the time-intensive matrix inversion processes. Methods such as Singular Value Decomposition (SVD), which decomposes a matrix into its singular values and vectors, and LU decomposition, which factors a matrix into a lower triangular matrix (L) and an upper triangular matrix (U), are much slower than the simple operations used in our approach.
For the rest of the paper, we abbreviate the actual FIM at $\theta^{k}$, ${\bm{F}}(\theta^{k})$ as ${\bm{F}}^{k}$ and the estimate of a batch as $\hat{{\bm{F}}}^{k}$ for notational convenience.
Sherman-Morrison for computing the inverse Fisher
$({\hat{\bm{F}}^{k}})^{-1}g^{k}$ $=\left[(\hat{\bm{F}}^{k-1}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\right]^{-1}g^{k}$
$\Delta\theta=\eta(\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top})^{-1}\log\pi_{\theta}(a|s),$
where $\lambda>0$ is a fixed damping coefficient.
@@ 4 NPG Approximation
We adopt a matrix-free scheme for approximating the FIM to reduce memory and computational cost. At each step, we form a local empirical Fisher from the current batch, add a damping term $\lambda$, and apply the Sherman–Morrison formula to compute the inverse–vector product ${\bm{F}}(\theta)^{-1}g^{k}$ directly, where $g^{k}$ is the policy gradient. This approach requires $\mathcal{O}(d)$ memory.
@@ 4.1 Matrix-Free Update
The time and space complexity of explicitly computing and storing the FIM is $\mathcal{O}(d^{2})$. So, maintaining a previous estimate $(\hat{{\bm{F}}}^{k-1})^{-1}$ and applying $(\hat{{\bm{F}}}^{k})^{-1}g$ can be infeasible for large-scale models. In this work, we instead use the Sherman–Morrison formula to directly compute the product of the inverse empirical Fisher and the policy gradient.
We consider a one-sample empirical Fisher with damping, take $(s_{k},a_{k})\sim d^{\pi_{\theta^{k}}}(s)\pi_{\theta^{k}}(a\mid s)$, that is, one state-action pair visited when rolling out $\pi_{\theta^{k}}$:
$\hat{\bm{F}}^{k}=\lambda\bm{I}+\bm{F}^{k}(\theta)=\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta^{k}}(a_{k}|s_{k})\nabla_{\theta}\log\pi_{\theta^{k}}(a_{k}|s_{k})^{\top}.$
Applying the Sherman–Morrison formula then yields:
$({\hat{\bm{F}}^{k}})^{-1}$ $=(\lambda\bm{I}+\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})\,\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})^{\top})^{-1}$
$=\frac{1}{\lambda}\bm{I}-\frac{\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})\,\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})^{\top}}{\lambda^{2}+\lambda\,\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})^{\top}\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})}$
$=\frac{1}{\lambda}\bm{I}-\frac{\bm{F}^{k}}{\lambda^{2}+\lambda\,\mathrm{Tr}(\bm{F}^{k})},$ (4)
Multiplying the inverse Fisher by the gradient gives the natural policy gradient update direction
$\Delta\theta_{k}=\eta\,({\hat{\bm{F}}^{k}})^{-1}g^{k}=\eta\left[\frac{1}{\lambda}g^{k}-\frac{\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})\,\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})^{\top}g^{k}}{\lambda^{2}+\lambda\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})^{\top}\,\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})}\right],$ (5)
where $g^{k}=\nabla_{\theta}\log\pi_{\theta^{k}}(a^{k}|s^{k})A(s,a)$ is the one-sample policy gradient update direction, $A(s,a)$ is the advantage estimate and $\eta$ is the step size.
@@ 4.2 Convergence Analysis
We take inspiration from the global convergence guarantees provided by Agarwal et al. (2021) and a follow-up work by Liu et al. (2022), which improved the global convergence results. To prove the global convergence, we take a standard Lipschitz policy assumption (Assumption 3) and average performance difference bounds (Lemma 3 from Liu et al. (2022)) and Lemma 2. We provide a theoretical analysis showing that the Sherman–Morrison rank-1 approximation to the inverse Fisher achieves global convergence. Under strong convexity assumptions (already taken in Kakade (2001); Peters et al. (2005)) of the FIM, we show that the average performance between the Sherman-Morrison update and the optimal policy $\pi^{*}$ can be bounded in terms of the the variance of policy gradient estimator ($\sigma^{2}$), discount factor ($\gamma$), the error in the Advantage estimate induced by the neural approximator ($\varepsilon_{bias}$) and the Lipschitz constant of the policy gradient ($L_{J}$). We defer the proof to the appendix, and state the final theorem here.
@@
In total, stochastic PG samples $\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{(1-\gamma)^{2}\varepsilon^{4}}\right)$ trajectories, where $\varepsilon_{\text{bias}}$ is the approximation error from the advantage function, and $\varepsilon$ is an arbitrary precision level and $\tilde{\mathcal{O}}$ hides all the logarithmic factors.
@@
Our proof relies on the fact that performance bound given by Lemma 3 (Liu et al., 2022) and we prove that this result holds with a sample complexity of $KN=\mathcal{O}(\frac{\sigma^{2}}{(1-\gamma)^{2}\varepsilon^{4}})$, where $N$ is the number of trajectories per update and $K$ is the number of updates. This shows that to get an $\varepsilon$ accurate policy, one needs $\mathcal{O}(\frac{1}{\varepsilon^{2}})$ samples, and when the horizon $H$ and number of samples $N$ are large enough, then $\sigma$ can be reduced to arbitrary precision. We would like to point out that the non-degeneracy of FIM [Assumption 2] and compatible function approximator [Assumption 4] are standard assumptions in stationary convergence of PG adapted by Agarwal et al. (2021) and can be realized by gaussian and softmax parametrized policies (See Section B.2 of Liu et al. (2022) and Section 8 of Ding et al. (2022) for further discussion). For compatible function approximation the value of $\varepsilon_{bias}$ can be arbitrarily small for neural approximators (Wang et al., 2019) and is exactly zero for softmax parametrized policies (Ding et al., 2022). Detailed derivations can be found in Appendix 2.
Note that, the obtained bound stems from average regret to the global optimum and recently Yuan et al. (2022) have obtained $\tilde{\mathcal{O}}(\varepsilon^{-3})$ bound on the condition when the approximator error is negligible, or $\varepsilon_{bias}=0$, but with a finite $\varepsilon_{bias}$ the authors have recovered a bound of $\tilde{\mathcal{O}}(\varepsilon^{-4})$, similar to ours.
4.3 A Toy Problem
x=0x=0x=1x=1a=0;r=+1\,a=0;\;r=+1\,a=0;r=+2\,a=0;\;r=+2\,a=1;r=0\,a=1;\;r=0\,a=1;r=0\,a=1;\;r=0\,
Figure 1: An example two-state MDP where each state $x$ has two control actions, one which self-transition with reward +1+1 or +2+2) and another control that transitions the initial state to the other state with zero reward).
We adopt a classical reinforcement–learning MDP problem, been used by Kakade (2001) & Bagnell and Schneider (2003) with two states $x\in\{0,1\}$ and two discrete actions $a\in\{0,1\}$. The policy is parameterized by a vector $\bm{\theta}=[\theta_{0},\theta_{1}]^{\top}$, where each component governs an independent Bernoulli distribution over actions. Let the probability of taking action $a=0$ in state $x$ be
$\pi_{\bm{\theta}}(a=0\mid x=0)$ $=\frac{1}{1+e^{\lambda_{p}\theta_{0}}},$ (6)
$\pi_{\bm{\theta}}(a=0\mid x=1)$ $=\frac{1}{1+e^{\theta_{1}}}$ (7)
where $\lambda_{p}$ scales the first policy parameter and is used to test scaling invariance. Assume that the agent starts at state $x=0$, then the objective function of the the MDP can be expressed as
$J(\bm{\theta})$ $=\pi_{\bm{\theta}}(a=0\mid x=0)+2(1-\pi_{\bm{\theta}}(a=0\mid x=0))\,\pi_{\bm{\theta}}(a=0\mid x=1).$ (8)
Note that, for a converged policy the agent shall transit to state $x=1$ and remain there and the final return shall be +2+2 which is the the action of self transitioning at state $x=1$. So we start from writing the analytic policy gradient vector,
$\nabla_{\bm{\theta}}J(\bm{\theta})=\begin{bmatrix}-\lambda_{p}\,\pi_{0}(1-\pi_{0})(1-2\pi_{1})\\[6.0pt] -2(1-\pi_{0})\pi_{1}(1-\pi_{1})\end{bmatrix}$. (9)
and the FIM can be calculated as:
$F(\bm{\theta})=\begin{bmatrix}\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})&0\\[4.0pt] 0&\pi_{1}(1-\pi_{1})\end{bmatrix}$. (10)
Then NPG update direction can be evaluated as
$\dot{\bm{\theta}}_{\text{NPG}}=F(\bm{\theta})^{-1}\,\nabla_{\bm{\theta}}J(\bm{\theta})=\begin{bmatrix}\dfrac{-(1-2\pi_{1})}{\lambda_{p}}\\[8.0pt] -2(1-\pi_{0})\end{bmatrix}$. (11)
This direction eliminates the dependence on the arbitrary scaling $\lambda_{p}$, making it invariant under smooth reparameterizations of $\bm{\theta}$ as shown by Bagnell and Schneider (2003).
Now the SM update direction in this setting can be evaluated as
$\dot{\bm{\theta}}_{\text{SM}}$ $=\frac{1}{\lambda}\left({\bm{I}}-\frac{F(\bm{\theta})}{\lambda+Tr(F(\bm{\theta}))}\right)\nabla_{\bm{\theta}}J(\bm{\theta})$ (12)
$=\begin{bmatrix}\dfrac{1}{\lambda}\!\left(-\lambda_{p}\,\pi_{0}(1-\pi_{0})(1-2\pi_{1})\right)-\dfrac{\lambda_{p}^{3}\,\pi_{0}^{2}(1-\pi_{0})^{2}(1-2\pi_{1})}{\lambda^{2}+\lambda\!\left[\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})+\pi_{1}(1-\pi_{1})\right]}\\ \dfrac{1}{\lambda}\!\left(-2(1-\pi_{0})\pi_{1}(1-\pi_{1})\right)-\dfrac{2\,\pi_{1}^{2}(1-\pi_{1})^{2}(1-\pi_{0})}{\lambda^{2}+\lambda\!\left[\lambda_{p}^{2}\,\pi_{0}(1-\pi_{0})+\pi_{1}(1-\pi_{1})\right]}\end{bmatrix}$ (13)
where $\lambda$ is the damping factor. This type of approximation although looks highly nonlinear but we can carefully tune $\lambda$ and make the update less sensitive to $\lambda_{p}$.
Figure 2: Results on toy MDP problem: Left side plot represents the log-odds plot for state $x=0$ in x-axis and state $x=1$ in y-axis and the right side plot shows the returns over iteration for PG, NPG and SM updates for different $\lambda_{p}\in\{0.75,1.0,2.0\}$ coefficient in policy parametrization. Note that, the least intense lines represent PG direction, medium is NPG and maximum deep lines represent SM update direction.
As shown in Fig.2, the plot in the left panel update the trajectories in the log-odds space for all damping settings $\lambda_{p}\in\{0.75,1.0,2.0\}$, Compared with the PG updates, the SM updates follow a direction much closer to that of NPG and maintain a step size more consistent. And PG updates deviate significantly as $\lambda$ changes, showing its sensitivity to parameter scaling. The plot in the right panel shows the returns over iterations, where SM update achieves faster and more stable improvement than PG and approaches the same final performance as NPG. Therefore, the SM update preserves the geometric characteristics of the NPG while being more computational efficient in high-dimensional cases.